A curve is represented by y = sin x. If x is changed from π/3 to π/3+π/100 , find approximately the change in y.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. y = sinx So, y + ∆y = sin (x + ∆x) ∆y = sin (x + ∆x) – sin x = (π/3+ π/100) – sin π/3 = 0.0157.